A high school reports that its students SAT scores were down by 20% for one year. The next year, however, SAT scores rose by 30%. The high school principal announces, question marks overall test scores have improved by 10% over the past two years.
1
Expert's answer
2019-09-27T13:50:24-0400

Let X be SAT scores. SAT scores were down by 20% for one year. It means it has become the value Y = (100% - 20%) from Y = 80% from X , or it has become Y = 0.8* X. So next year SAT scores rose by 30%. It means it has become (100% + 30 %) from Y = 130% from Y = 1.3* Y = 1.3* ( 0.8* X ) = 1.04 * X = 100 % + 4% from X. So it means SAT scores has improved by 4 % over the past two years. The high school principal was wrong to say it was 10%.